OpenStudy (anonymous):

What is the equation of the line which includes points (1, 1) and (1, 5)?

OpenStudy (jdoe0001):

\(\large \bf {\begin{array}{lllll} &x_1&y_1&x_2&y_2\\ &(1\quad ,&1)\quad &(1\quad ,&5) \end{array} \\\quad \\ slope = m= \cfrac{rise}{run} \implies \cfrac{y_2-y_1}{x_2-x_1} \\ \quad \\ y-y_1=m(x-x_1)\quad \textit{plug in your values and solve for "y"}}\)